OpenStudy (anonymous):

Find the volume of a cylinder that has a diameter of 8 cm and a height of 15 cm. can you do it and sho all of your teps It is over due and I need to turn it in fast I will fan and give a medal, thanks!!!

OpenStudy (leilanilane):

Easy! Okay: Formula - V = pi * r^2 * h Pi - 3.14 r - 8 h - 15 8 * 8 = 64 64 * 15 = 960 960 * 3.14 = 3014.4 cm

OpenStudy (leilanilane):

Do you get it? O.O

OpenStudy (anonymous):

Thank you so much thanks for you I might not end up working at Mcdonalds;) lol haha

OpenStudy (leilanilane):

Haha :3 I've actually got some math homework of my own to start working on :S If you need anymore help, don't hesitate to ask! ;D